What is the volume of 2.43 x 10^23 molecules of N2 gas at STP.


1
Expert's answer
2021-05-08T23:59:19-0400

6.02×1023 molecules = 1 mole 

2.43×1023 molecules = 0.404 mole 

But PV = nRT 

where P = pressure of gas , V = volume 

n = no. of moles , T = Temperature, R = gas constant 

P=1atm

R=0.0821L/atm/mol/K

T=273K at STP

n=0.404moles


But V="\frac{nRT}{P}"


="\frac{0.404 mole ×0.0821L/atm/mol/K×273K}{1atm}"


V=9.05L